it takes Dave 12 hours to clean the kitchen at his restaurant after throwing a legit birthday party for his buddy, Buster. When Dave and Buster clean the kitchen together, they can get it done in 5 and a half hours. How long does it take Dave to clean the kitchen by himself (when it isn’t his birthday)?


Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

Dave can clean 1/12 of the kitchen in an hour.

1/12 + 1/x = 11/2
x = 12/65
It takes Buster 12/65 hours.

Answer by ikleyn(52803)

.
it takes Dave 12 hours to clean the kitchen at his restaurant after throwing a legit birthday party for his buddy, Buster.
When Dave and Buster clean the kitchen together, they can get it done in 5 and a half hours.
How long does it take Dave to clean the kitchen by himself (when it isn’t his birthday)?
~~~~~~~~~~~~~~


Hello,


            in the first sentence of the post, the problem says "it takes  Dave  12  hours to clean the kitchen".

            Then, at the end, the post asks again  "How long does it take  Dave to clean the kitchen by himself ?".


The answer is  "12 hours",  as it is stated in the first statement.


My impression is that the author  (the problem's composer)  was so infatuate by his writing
that missed Dave and Buster each with the other somewhere in between the beginning and the end of the post.


Double check your post.


If you will find it necessary to re-write the post, then do it and re-post to the forum.